14. + 0/1 points Previous Answers OSUniPhys1 13.4.P.040. My Notes Ask Your Teacher For the following questions, use the dataCouldn't get the answer no matter what...Please make sure you get it right, so I can give you a thumbs up!
Answer #1

T = 29.4*31.53610^6 = 927.1584*10^6 S

G = 6.67*10^-11 N.m^2/kg^2

r = 143.4*10^10 m

M sun = 4*pi^2 *(143.4*10^10)^3/(6.67*10^-11*(927.1584*10^6)^2

M sun = 2.0303*10^30 kg (calculated )

M sun = 1.989*10^30 kg (actual )

M calculated / Mactual = 2.0303*10^30 kg / 1.989*10^30 kg =

M calculated / Mactual = 1.0207